# Wavy Transformer

### v1

@@ Appendix A Energy Evolution in Diffusion and Wave Systems
$\frac{d}{dt}\int{\left(x(u,t)\right)^{2}}du=-2\int|\nabla x(u,t)|^{2}du\leq 0,$
@@
Define the energy functional for $x=x(u,t)$ which is a solution to the wave equation
$E(t)=\frac{1}{2}\int\left[\left(\frac{\partial{x(u,t)}}{\partial t}\right)^{2}+|\nabla x(u,t)|^{2}\right]du.$
@@
Appendix B Appendices of Difference Between Pre-LN and Post-LN
@@
B.2 Pre-LN vs. Post-LN Implementations in Wavy Transformer
@@
For considering Wavy Transformer,the difference between Pre-LN and Post-LN xiong2020layer is not trivial. Therefore, this section examines the implementation differences between Pre-LN and Post-LN in the context of Wavy Transformer. Building on the Pre-LN and Post-LN baseline implementations shown Table 4, we construct Post-LN and Pre-LN versions of Wavy Transformer, whose details are summarized in Table 5. Conceptually, Wavy Transformer replaces the conventional residual update based on diffusive dynamics with one governed by wavy dynamics.
@@
B.3 Attention with Pre-LN as graph neural diffusion with reaction term
@@
$\frac{\partial x(u,t)}{\partial t}=\underbrace{\mathrm{div}\bigl{(}D\,\nabla x(u,t)\bigr{)}}_{\text{diffusion}}+\underbrace{b(u,t)}_{\text{source}}-\underbrace{\kappa(u,t)\,x}_{\text{reaction}}.$ (23)
@@
Appendix C Intuitive Explanation of Over-smoothing
@@
Appendix D Analysis of Over-smoothing for DeiT-based models
(a) Comparison between cosine similarities with 1-sigma interval shading of the conventional DeiT and the DeiT integrated with Wavy Transformer.
(b) Comparison between cosine similarities with 1-sigma interval shading of the DeiT integrated with Wavy Transformer with and without FeatScale.
In this section, we analyze over-smoothing in DeiT-based models on the ImageNet classification task. Following the NLP approach, we quantify over-smoothing via the average pairwise cosine similarity between patch tokens, computed on the ImageNet validation dataset. Figure 5 plots these similarity curves for the standard DeiT, DeiT integrated with Wavy Transformer, and Wavy Transformer + FeatScale variant with 1-sigma interval shading. The results clearly show how Wavy Transformer reshapes feature dynamics relative to the conventional DeiT. Integrating FeatScale slightly reduces over-smoothing overall, although its final-layer cosine similarity remains marginally higher than that of Wavy Transformer alone.
Appendix E Energy Evaluation in Diffusion and Wave Dynamics on Graphs
In this section, to indicate an intuitive discussion of diffusion and wave dynamics on graphs from an energetic viewpoint, we assume here time-independent attention matrix $\mathbf{A}$ (e.g. $\mathbf{A}_{ij}=\mathbf{A}(\mathbf{X}^{0}_{i},\mathbf{X}^{0}_{j})$ using the initial features) and impose the symmetry condition $\mathbf{A}_{ij}=\mathbf{A}_{ji}$. These simplifying assumptions do not hold exactly in real transformers; they are introduced here solely for analytical simplicity and an intuitive understanding of their hidden-state dynamics. Note that this section presents the discrete-graph analogue of Appendix A and should be compared with it.
@@ Appendix E Energy Evaluation in Diffusion and Wave Dynamics on Graphs
This shows that the hidden-state dynamics in transformers, which implicitly rely on graph neural diffusion, can be interpreted as the gradient flow of the potential energy in Eq. (28). A similar discussion appears in nguyen2023mitigating ; 10.5555/3666122.3667319 , and a more general mathematical discussion is given in geshkovski2023mathematical .
Let the node features $\mathbf{X}$ evolve according to the graph neural diffusion in Eq. (7). To clarify the meaning of the potential energy, we use the symmetry of $\mathbf{A}$ to rewrite it as
@@
where $\|\mathbf{X}_{j}-\mathbf{X}_{i}\|^{2}=(\mathbf{X}_{j}-\mathbf{X}_{i})^{\top}(\mathbf{X}_{j}-\mathbf{X}_{i})$. Taking the time derivative gives
$\frac{dU(\mathbf{X})}{dt}$ $=\frac{1}{4}\sum_{i,j}\mathbf{A}_{ij}\,\frac{d}{dt}\bigl{\|}\mathbf{X}_{j}-\mathbf{X}_{i}\bigr{\|}^{2}$ (33)
$=\frac{1}{2}\sum_{i,j}\mathbf{A}_{ij}\,\bigl{(}\dot{\mathbf{X}}_{j}-\dot{\mathbf{X}}_{i}\bigr{)}^{\top}\bigl{(}\mathbf{X}_{j}-\mathbf{X}_{i}\bigr{)}$ (34)
$=-\frac{1}{2}\sum_{i,j}\bigl{\|}\mathbf{A}_{ij}(\mathbf{X}_{j}-\mathbf{X}_{i})\bigr{\|}^{2}\;\leq\;0,$ (35)
where
$\dot{\mathbf{X}}_{j}-\dot{\mathbf{X}}_{i}=\bigl{(}\mathbf{A}_{ji}\mathbf{X}_{i}-\mathbf{X}_{j}\bigr{)}-\bigl{(}\mathbf{A}_{ij}\mathbf{X}_{j}-\mathbf{X}_{i}\bigr{)}.$ (36)
If we define the attention-weighted average $\bar{\mathbf{X}}=\mathbf{A}\mathbf{X}$ (cf. Appendix C), then
$\frac{dU(\mathbf{X})}{dt}=-\frac{1}{2}\sum_{i}\bigl{\|}\bar{\mathbf{X}}_{i}-\mathbf{X}_{i}\bigr{\|}^{2}.$ (37)
This demonstrates that the graph neural diffusion drives the node features $\mathbf{X}$ to decrease the potential energy defined in Eq. (32), and that $U(\mathbf{X})$ can be interpreted as the sum of squared differences between each feature and its attention-weighted average, implying convergence toward a uniform state. This provides a clear explanation for the over-smoothing observed in transformers from the energetic viewpoint.
@@
Appendix F Throughput–Accuracy Trade-off Comparison
Table 6 compares the inference throughput (images/s) of various transformer variants, measured on the ImageNet validation set using V100 GPUs and a batch size of 256. Introducing wave dynamics (Diffuse+Wave or Wave alone) reduces throughput by roughly 50% from about 2 600 to 1 200–1 300 images per second which may pose a limitation in real-time applications. However, this slowdown must be balanced against the gains in representational power and accuracy that wave-enhanced models provide. A simple mitigation is to insert wavy blocks in only a subset of layers to recover part of the lost throughput. Concretely, replacing only the last six blocks with wavy blocks (Wave (6)) increases throughput by about 33 % relative to the full Diffuse+Wave with FeatScale variant (1 649 img/s vs. 1 241 img/s) while sacrificing 0.18 pt in Top-1 accuracy (72.44 % vs. 72.62 %). In this experiment, the velocity tensor was initialized to zero at the first wavy block.17
Model Residual Connections Throughput Top-1 (%)
DeiT-Ti DeiT Diffuse 2632.1 72.17
DeiT-Ti + FeatScale wang2022antioversmooth Diffuse 2483.2 72.35
DeiT-Ti Diffuse+Wave 1266.0 72.33
DeiT-Ti + FeatScale Diffuse+Wave 1241.4 72.62
DeiT-Ti Wave 1312.2
DeiT-Ti + FeatScale Diffuse+Wave (6) 1649.3 72.44
Appendix G Implementation Details
Table 7: Hyper-parameters for different downstream tasks.
GLUE SQuAD
Batch size 32 16
Weight decay [0.1, 0.01] [0.1, 0.01]
Warmup proportion 0.1 0.1
Learning rate decay linear linear
Training epochs 3 3
Learning rate 0.00005 0.00005

### v2

@@ Appendix C Energy Evolution in Diffusion and Wave Systems
$\frac{d}{dt}\int{\left(x(s,t)\right)^{2}}ds=-2\int|\nabla x(s,t)|^{2}ds\leq 0,$
@@
Define the energy functional for $x=x(s,t)$ which is a solution to the wave equation
$E(t)=\frac{1}{2}\int\left[\left(\frac{\partial{x(s,t)}}{\partial t}\right)^{2}+|\nabla x(s,t)|^{2}\right]ds.$
@@
Appendix D Appendices of Difference Between Pre-LN and Post-LN
@@
D.2 Pre-LN vs. Post-LN Implementations in Wavy Transformer
@@
For considering Wavy Transformer with Full Wave setting, the difference between Pre-LN and Post-LN xiong2020layer is not trivial. Therefore, this section examines the implementation differences between Pre-LN and Post-LN in the context of Wavy Transformer. Building on the Pre-LN and Post-LN baseline implementations shown Table 18, we construct Post-LN and Pre-LN versions of Wavy Transformer, whose details are summarized in Table 19. Conceptually, Wavy Transformer replaces the conventional residual update based on diffusive dynamics with one governed by wavy dynamics.
@@
D.3 Attention with Pre-LN as graph neural diffusion with reaction term
@@
$\frac{\partial x(s,t)}{\partial t}=\underbrace{\mathrm{div}\bigl(D\,\nabla x(s,t)\bigr)}_{\text{diffusion}}+\underbrace{b(u,t)}_{\text{source}}-\underbrace{\kappa(u,t)\,x}_{\text{reaction}}.$ (27)
@@
Appendix E Intuitive Explanation of Over-smoothing
@@
Appendix F Energy Evaluation in Diffusion and Wave Dynamics on Graphs
In this section, to indicate an intuitive discussion of diffusion and wave dynamics on graphs from an energetic viewpoint, we assume here time-independent attention matrix $\mathbf{A}$ (e.g. $\mathbf{A}_{ij}=\mathbf{A}(\mathbf{X}^{0}_{i},\mathbf{X}^{0}_{j})$ using the initial features) and impose the symmetry condition $\mathbf{A}_{ij}=\mathbf{A}_{ji}$. These simplifying assumptions do not hold exactly in real transformers; they are introduced here solely for analytical simplicity and an intuitive understanding of their hidden‐state dynamics. Note that this section presents the discrete‐graph analogue of Appendix C and should be compared with it.
@@ Appendix F Energy Evaluation in Diffusion and Wave Dynamics on Graphs
This shows that the hidden‐state dynamics in transformers, which implicitly rely on graph neural diffusion, can be interpreted as the gradient flow of the potential energy in Eq. (32). A similar discussion appears in nguyen2023mitigating ; 10.5555/3666122.3667319 , and a more general mathematical discussion is given in geshkovski2023mathematical .
Let the node features $\mathbf{X}$ evolve according to the graph neural diffusion in Eq. (7). Using the symmetry of $\mathbf{A}$ and the row-stochastic condition $\sum_{j}\mathbf{A}_{ij}=1$, we can rewrite the potential as
@@
where $\|\mathbf{X}_{j}-\mathbf{X}_{i}\|^{2}=(\mathbf{X}_{j}-\mathbf{X}_{i})^{\top}(\mathbf{X}_{j}-\mathbf{X}_{i})$. Taking the time derivative along the gradient flow yields
$\frac{dU(\mathbf{X})}{dt}$ $=\biggl\langle\frac{\partial U(\mathbf{X})}{\partial\mathbf{X}},\,\frac{\partial\mathbf{X}}{\partial t}\biggr\rangle\;=\;-\biggl\|\frac{\partial U(\mathbf{X})}{\partial\mathbf{X}}\biggr\|^{2}\;=\;-\bigl\|(\mathbf{I}-\mathbf{A})\,\mathbf{X}\bigr\|^{2}\;\leq\;0,$ (37)
where $\langle\cdot,\cdot\rangle$ denotes the Frobenius inner product, i.e., $\langle\mathbf{Y},\mathbf{Z}\rangle=\mathrm{tr}(\mathbf{Y}^{\top}\mathbf{Z})$. If we define the attention‐weighted average $\bar{\mathbf{X}}=\mathbf{A}\mathbf{X}$ (cf. Appendix E), then
$\frac{dU(\mathbf{X})}{dt}\;=\;-\,\|\bar{\mathbf{X}}-\mathbf{X}\|^{2}\;=\;-\sum_{i}\bigl\|\bar{\mathbf{X}}_{i}-\mathbf{X}_{i}\bigr\|^{2}.$ (38)
This demonstrates that the graph neural diffusion decreases the potential energy (36) monotonically, and that $U(\mathbf{X})$ can be interpreted as the sum of squared differences between each feature and its attention‐weighted average, implying convergence toward a uniform state. This provides a clear explanation for the over‐smoothing observed in transformers from the energetic viewpoint.